OpenStudy (anonymous):

How do I find the 4th degree Taylor polynomial for cos2(x).

OpenStudy (anonymous):

cos^2(x)

OpenStudy (anonymous):

Just find the Taylor Polynomial up to fourth term expanding the formula \[\sum_{k=0}^{\infty}(f ^{k}(c).(x-c)^{k})\div k!\] Note that you have been given the value of c (like c=2 or c=0...) in order to find the Taylor Polynomial.

OpenStudy (anonymous):

If you are confused in finding the derivative \[\cos ^{2}x = (1 + \cos 2x) / 2\] and find the derivative.